OpenStudy (anonymous):

1/(b^{-2}+a^{-2})= b^{2}+a^{2} True or False?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

left hand side is \[\frac{1}{\frac{1}{b^2}+\frac{1}{a^2}}\]

OpenStudy (anonymous):

ohh that makes sense is it mainly because their being added?

OpenStudy (anonymous):

you could get rid of this complex fraction by multiplying top and bottom by \(a^2b^2\)

OpenStudy (anonymous):

you would get \[\frac{a^2b^2}{a^2+b^2}\] which is certainly not \(a^2+b^2\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks satellite 73

OpenStudy (anonymous):

yw